OpenStudy (anonymous):

Fourier Transform of f(t) = t h(t), where h(t) is the heaviside step function.

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}t e ^{iwt}\]

OpenStudy (anonymous):

But how do I solve it? Im not quite sure how to do it by hand

OpenStudy (anonymous):

What's your definition for the step function?

OpenStudy (anonymous):

The way we've been using it is 0 for t < 0, 1 for t > 0. Sorry for the late response

OpenStudy (anonymous):

Okay, so it's the standard definition. Your integral setup is correct. Try integrating by parts.

OpenStudy (anonymous):

Where \[u = t, du = dt; dv = e ^{iwt} dt, v = \frac{ 1 }{ iw } e ^{iwt}\] \[\left[ \frac{ t }{ iw }e ^{iwt} \right] 0 \to \infty - \int\limits_{0}^{\infty} \frac{ 1 }{ wt }e ^{iwt}dt\]

OpenStudy (anonymous):

I feel like doing it the other way made it worse. These integrals go to infinity o_0

OpenStudy (anonymous):

If I recall correctly, there's a formula for Fourier transforms that says \[\large\mathcal{F}\left\{t~u(t)~e^{-\alpha t}\right\}=\frac{1}{(\alpha+i\omega)^2}\] In this case, you'd have \(\alpha=0\), so it looks like that integral should converge...

OpenStudy (anonymous):

In mathematica, I got something in terms of a delta'(w), one sec...

OpenStudy (anonymous):

Yes I'm getting the same result, \[-\frac{1}{\sqrt{2\pi}\omega^2}-i\sqrt\frac{\pi}{2}\delta'(\omega)\]

OpenStudy (anonymous):

Yet integrating manually (via Mma, at least) yields \(-\dfrac{1}{\omega^2}\), which agrees with that formula.

OpenStudy (anonymous):

Can you explain that part please? I wasnt sure how to get the complicated result but I see now how to get the result from the transform formula you showed me just now

OpenStudy (anonymous):

I'm trying to figure out the difference myself ;)

OpenStudy (anonymous):

Haha, alright, thanks for the help again

OpenStudy (anonymous):

And where can I find that transform formula/others like it by chance?

OpenStudy (anonymous):

Here's one table I just found (bottom of third page): http://uspas.fnal.gov/materials/11ODU/FourierTransformPairs.pdf

OpenStudy (anonymous):

I think the discrepancy is that Mathematica uses a different definition for the transform.

OpenStudy (anonymous):

Ah, just missed that one. Thank you

OpenStudy (anonymous):

Yeah, it uses the different coefficient =/

OpenStudy (anonymous):

What Mma uses: http://mathworld.wolfram.com/FourierTransform.html What you seem to be using: \[\large\mathcal{F}\{f(t)\}:=\int_{-\infty}^\infty f(t)e^{-i\omega t}~dt\]

OpenStudy (anonymous):

I think, in your work above, you're missing the negative in the exponent. This should fix the result.

OpenStudy (anonymous):

I blame the teacher. He has it opposite and I have been trying to figure out why. He uses positive for t-> frequencey and negative for backwards. Then he uses negative/positive for x -> k

OpenStudy (anonymous):

Why does the MM answer equal the simpler answer?

OpenStudy (anonymous):

Im not sure what to do with delta prime... or how to understand it anyways

OpenStudy (anonymous):

I think the software uses a different algorithm for `FourierTransform[]` than it does for `Integrate[]`.

OpenStudy (anonymous):

It might have to do with its definition of the step function, which seems to be zero for \(t<0\) and 1 for \(t>0\) as opposed to \(\le0\) or \(\ge1\).

OpenStudy (anonymous):

Gotcha, ill try to look into some more. Thanks for everything!

OpenStudy (anonymous):

Wait. Does the above Fourier Transform formula work if e isnt explicitly in the function?

OpenStudy (anonymous):

f(t) = t h(t)

OpenStudy (anonymous):

You can always write any function that doesn't explicitly contain an exponential function by tacking on a factor of \(e^0\).

OpenStudy (anonymous):

True point

OpenStudy (anonymous):

So what we could have done was multiply the transform of t by the transform of u(t) ? This gives me....